OpenStudy (anonymous):

Find the derivative of f(x) = negative 7 divided by x at x = -3.

OpenStudy (anonymous):

The derivative:\[\frac{ d }{ dx }(-\frac{ 7 }{ x })=\frac{ 7 }{ x^2 }\] Now insert \[x=-3\]

OpenStudy (anonymous):

7/9